# FIRM: Federated In-client Regularized Multi-objective Alignment for Large Language Models

### v1

@@ 1 Introduction
Large Language Models (LLMs) have become indispensable in applications ranging from digital assistants to scientific discovery (snell2023offline; pyatkin2023clarifydelphi). Yet their deployment is inseparable from the question of alignment: ensuring that model behavior is consistent with human values such as helpfulness and harmlessness (ouyang2022training). Crucially, alignment is not a single-objective problem but a balancing act across multiple, often conflicting, goals. For example, maximizing helpfulness can come at the expense of safety, while enforcing strict harmlessness may lead to evasive or uninformative answers. Designing scalable methods to navigate such trade-offs has therefore emerged as a central challenge in the responsible development of LLMs.
@@
The dominant alignment paradigm, Reinforcement Learning from Human Feedback (RLHF) and its recent variants like Direct Preference Optimization (DPO) and Group Relative Policy Optimization (GRPO) (rafailov2023direct; shao2024deepseekmath), optimizes policies from preference data. While successful, these methods are fundamentally centralized, demanding vast labeled datasets and large-scale, proprietary infrastructure. This centralization creates significant barriers to entry, limiting broader research participation and raising pressing data privacy concerns. Federated Learning (FL) offers a compelling alternative (mcmahan2017communication). By training models locally on client data and aggregating only model updates, FL presents a path toward democratizing alignment research while inherently preserving privacy. However, applying FL to multi-objective LLM alignment introduces a new, fundamental design question: Should the resolution of conflicting objectives happen at the central server or locally at each client?
@@
Recent work in federated multi-objective supervised learning has explored the server-centric approach to conflict resolution. A straightforward implementation requires each client to transmit gradients for all $M$ objectives to a central server, which then finds a globally consistent update direction (yang2023federated). This design is practically untenable for modern LLMs, as it inflates communication costs by a factor of $M$, resulting in a prohibitive overhead of $\mathcal{O}(CMd)$, where $C$ denotes the number of clients and $d$ the model dimension. Recognizing this bottleneck, other works have proposed using gradient compression before transmission (askin2024federated). However, this remedy introduces significant new overheads: convergence becomes fundamentally limited by the quality of the compression, an extra communication round-trip is required in every step, and the compression itself can be computationally prohibitive for models with billions of parameters. Thus, whether naive or compressed, the server-centric paradigm remains a non-starter for efficient, large-scale LLM alignment.
@@
As our first contribution, we shift the paradigm to a communication-efficient client-centric approach. We introduce Federated In-client Regularized Multi-objective alignment (FIRM), the first framework, to our knowledge, for federated multi-objective LLM alignment. In this framework, each client computes gradients for all $M$ objectives using Proximal Policy Optimization (PPO) (schulman2017proximal), and then resolves them locally into a single update direction by solving a Multiple-Gradient Descent Algorithm (MGDA) subproblem (desideri2012multiple). The central server’s role is simply to aggregate the updated client parameters via FedAvg (mcmahan2017communication), which reduces communication costs to a practical $\mathcal{O}(Cd)$. However, in deploying this method, we identified a critical instability inherent to naive client-side MGDA resolution, which we term multi-objective disagreement drift. Because the solution to the local MGDA subproblem is highly sensitive to input gradients, minor stochastic variations across clients cause their computed update directions to diverge significantly. When aggregated at the server, these conflicting updates impede learning and prevent effective alignment (see Section 3 and Remark 3 in Section 4.2).
@@ 2 Problem Formulation
Finding Pareto-optimal policies in non-convex settings such as LLM alignment is NP-hard. Thus, we instead target Pareto stationarity, a first-order necessary condition for Pareto optimality (desideri2012multiple; zhou2024finite).
@@ 3 The FIRM Algorithm for LLM Alignment
In this section, we present our proposed algorithm, Federated In-client Regularized Multi-objective alignment (FIRM). FIRM is designed for practical LLM alignment and proceeds in communication rounds, with each client performing multiple local updates before aggregation. The underlying RL algorithm is a multi-objective adaptation of Proximal Policy Optimization (PPO) (schulman2017proximal; zhou2024finite), a standard for LLM fine-tuning. For efficiency, clients fine-tune the LLM using LoRA, meaning only the low-rank adapter matrices are trained and communicated.
@@
We now explain why both terms $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are needed in Eq. (1). The first term, $\mathcal{T}_{1}$, corresponds to the MGDA (desideri2012multiple). MGDA seeks a convex combination of objective gradients that defines a common descent direction. In a federated setup, however, using only $\mathcal{T}_{1}$ is not sufficient. Without $\mathcal{T}_{2}$, client updates can drift apart. To see this, consider rewriting Eq. (1) as
@@ 4.1 Theoretical Setup and Assumptions
To enable a tractable convergence analysis, we analyze a variant of FIRM, which we refer to as Theoretical-FIRM (TFIRM), where the PPO update is replaced by a foundational actor-critic (AC) framework. This substitution allows us to isolate the core federated and multi-objective challenges. In this AC setting, each client’s local update proceeds in two steps. First, in an inner loop, the critic employs linear function approximation to update its $M$ value function estimates, $\{V_{w}^{j}\}_{j=1}^{M}$, via TD learning on mini-batches of trajectory data. Subsequently, the actor uses these estimates to form the TD-error as an advantage approximation for each objective $j\in[M]$, and then computes the per-objective gradients using the policy gradient theorem (sutton1999policy). Beyond local gradient computation, the federated optimization protocol, including server aggregation and our regularized MGDA solver, remains identical to Algorithm 1. Our theoretical framework is a novel extension of the multi-objective AC algorithm of zhou2024finite to the federated setting, with the regularized MGDA solver ensuring robust consensus on the descent direction. The complete algorithm is provided in Appendix C (See Algorithm 2), while in the remainder of this section we carry out our analysis under the following standard assumptions.
@@
Assumption 1 imposes standard regularity conditions (zhou2024finite). Condition (a) requires differentiability, which is essential for applying policy gradient methods. Condition (b) ensures ergodicity of the Markov process, yielding a unique stationary distribution under any policy and thereby guaranteeing that long-term objectives are well defined. Finally, condition (c) assumes bounded rewards, a standard requirement to ensure that the value functions remain bounded.
@@
Assumption 2 is standard in the analysis of linear temporal-difference learning (tsitsiklis1999average; xu2020improving; xu2020reanalysis; qiu2021finite; zhou2024finite). The conditions collectively ensure that the projected Bellman equation is well-posed for any policy $\pi_{\theta}$. Specifically, the negative definiteness of the matrix $\mathbf{A}_{\pi_{\theta}}$, combined with bounded features and rewards, guarantees that the optimal weight vector $\mathbf{w}_{\theta}^{*,j}$ that approximates the value function is unique and uniformly bounded across all policies.
@@
This assumption outlines standard smoothness and boundedness conditions essential for our analysis. The bounded score function (a), a property of common policies like softmax, prevents arbitrarily large gradient updates. The Lipschitz condition on the objective gradients (b) ensures a sufficiently smooth optimization landscape, which is a cornerstone for guaranteeing the convergence of gradient-ascent algorithms to a stationary point (xu2020improving).
@@ 4.2 Results and convergence analysis
Theorem 1 provides the first convergence guarantee for a federated multi-objective actor-critic algorithm, ensuring that TFIRM converges to a Pareto stationary point. The overall convergence rate, dominated by the optimization error, is $\tilde{\mathcal{O}}(1/T)$. This matches the rate for centralized multi-objective actor-critic methods (zhou2024finite), demonstrating that our federated approach achieves a comparable asymptotic performance. Furthermore, the variance term, $\mathcal{O}(1/(CB))$, exhibits the expected linear speedup with respect to the number of clients $C$ and batch size $B$, consistent with standard analyses in federated learning (zhang2024finitetime).
@@
The most significant insight from our analysis lies in the characterization of the client drift. The total drift consists of two components. The first, $\mathcal{O}(\alpha^{2}K^{2})$, is the classical drift from local updates, which is well-understood in federated optimization (zhou2024finite; zhang2024finitetime). The second, and the core of our theoretical contribution, is a novel error term we identify as the multi-objective disagreement drift, given by $\mathcal{O}\!\left(\frac{\sqrt{M^{3}}}{\beta\sqrt{B}}\alpha K\right)$. This term precisely captures the error arising from clients solving the MGDA problem locally with stochastic gradients. Our bound reveals two key insights: (i) the challenge of reaching consensus grows with the number of objectives $M$, and (ii) this drift can be explicitly controlled by the product of the regularization $\beta$ and the batch size $B$.
Comparing to askin2024federated
While our work is the first in the RL setting, the most relevant methodological comparison is with the federated multi-objective supervised learning framework of askin2024federated. Their server-centric approach tackles client disagreement by having clients communicate compressed gradients to the server, which then solves a single MGDA problem and broadcasts the solution vector $\lambda$ back. This design for supervised learning introduces an error term of $\mathcal{O}(qM)$ into the convergence upper bound, where the factor $q$ depends on the quality of the gradient compression performed at each client. This approach, therefore, introduces significant overhead. A new error source: the convergence is fundamentally limited by the quality of the gradient compression, as represented by the $q$ term. Communication cost: it requires an extra round-trip communication in every training step (gradients up, $\lambda$ down). Computational cost: it relies on a dimensionality reduction step on the gradients at each client, which can be computationally prohibitive for large models. In contrast, our approach removes the need for costly server-side coordination and its associated trade-offs. By introducing a lightweight $\ell_{2}$ regularization to each client’s local MGDA objective, we provably control the multi-objective disagreement drift.
@@
Our upper bound differs from askin2024federated by an additional $\sqrt{M}$ factor. This may stem from the fact that client drift in RL is more severe than in supervised learning, as local policy updates affect both the local model parameters and the local induced data distribution. Understanding whether this dependence is tight remains an interesting open problem left for future work.
@@
The critic approximation error $\varepsilon_{\text{critic}}$ in our bound can be controlled using standard results from temporal-difference learning. In particular, analyses of Markovian mini-batch TD show that this error can be driven arbitrarily close to zero with sufficient data. From Theorem 1 of xu2020improving, we obtain $\mathbb{E}[\|w-w^{*}\|^{2}]\leq\epsilon$ with sample complexity $\mathcal{O}(\epsilon^{-1}\log(1/\epsilon))$. Thus, in our convergence analysis, the critic error can be treated as a controllable parameter (see Lemma 3 in Appendix D).
@@ 5 Numerical Experiments
We fine-tune TinyLlama-1.1B-Chat-v1.0  [^1] The model choice reflects realistic edge-device constraints in federated learning, where resources are limited. (tinyllama) using LoRA adapters, ensuring only adapter weights are trained and communicated. Our alignment task uses $M=2$ objectives, Helpfulness and Harmlessness, with prompts from the Anthropic HH-RLHF (bai2022training) dataset and rewards evaluated by public reward models [^2] Ray2333/gpt2-large-helpful-reward_model (helpfulness) and Ray2333/gpt2-large-harmless-reward_model (harmlessness) (yang2024rewards). Moreover, all reward scores are normalized to the $[0,1]$ range. Each generated response receives a 2-D reward vector $(r_{\mathrm{help}},r_{\mathrm{harm}})$ which drives the PPO updates. Our federated protocol involves $C$ clients over $T$ rounds, where each client performs a fixed number of local PPO-style updates before the server aggregates the LoRA (rank r=16r{=}16 , $\alpha_{\mathrm{lora}}{=}32$, dropout $0.05$) adapters via FedAvg. For the local MGDA subproblem, we solve a regularized Quadratic Program (QP). However, the scale of the gradients, and thus their Gram matrix $G$, can vary dramatically during training. To ensure our regularization $\beta$ has a consistent effect, we first normalize $G$ by its trace, which stabilizes the local optimization (See Appendix A). We set the regularization parameter to $\beta=0.01$. All experiments were conducted on a server equipped with NVIDIA H100 GPUs.
@@
To evaluate FIRM, we compare it against a strong server-centric baseline, which we term FedCMOO-A. This baseline is our adaptation of the federated multi-objective supervised learning algorithm, FedCMOO (askin2024federated), to the LLM alignment setting. In FedCMOO-A, clients send their local multi-objective gradients to the server. The server then solves a single MGDA problem to compute a global trade-off vector $\lambda$ and broadcasts it back to the clients. This architecture avoids the multi-objective disagreement drift by design, as all clients are forced to use the same $\lambda$. To ensure a fair comparison focused purely on the conflict resolution strategy, we do not use gradient compression in FedCMOO (askin2024federated), thereby removing any potential compression error. As shown in Figure 1, FIRM, with 8 clients, achieves comparable or superior performance to FedCMOO-A. One hypothesis for FIRM’s strong performance is the agility of its local updates. FIRM clients can adjust their trade-off vectors $\lambda_{c}$ at every local step, allowing for a more responsive and potentially better balance between objectives. In contrast, FedCMOO-A clients must wait for the server’s global $\lambda$, which can become “stale” or lagged between communication rounds. This lag results in oscillatory weight trajectories, as the server repeatedly overcorrects its global $\lambda$ (See Figure 1(c) and Figure 1(d)).

### v2

@@ 1 Introduction
Large Language Models (LLMs) have become indispensable in applications ranging from digital assistants to scientific discovery (Snell et al., 2023; Pyatkin et al., 2023). Yet their deployment is inseparable from the question of alignment: ensuring that model behavior is consistent with human values such as helpfulness and harmlessness (Ouyang et al., 2022). Crucially, alignment is not a single-objective problem but a balancing act across multiple, often conflicting, goals. For example, maximizing helpfulness can come at the expense of safety, while enforcing strict harmlessness may lead to evasive or uninformative answers. Designing scalable methods to navigate such trade-offs has therefore emerged as a central challenge in the responsible development of LLMs.
@@
The dominant alignment paradigm, Reinforcement Learning from Human Feedback (RLHF) and its recent variants like Direct Preference Optimization (DPO) and Group Relative Policy Optimization (GRPO) (Rafailov et al., 2023; Shao et al., 2024), optimizes policies from preference data. While successful, these methods are fundamentally centralized, demanding vast labeled datasets and large-scale, proprietary infrastructure. This centralization creates significant barriers to entry, limiting broader research participation and raising pressing data privacy concerns. Federated Learning (FL) offers a compelling alternative (McMahan et al., 2017). By training models locally on client data and aggregating only model updates, FL presents a path toward democratizing alignment research while inherently preserving privacy. However, applying FL to multi-objective LLM alignment introduces a new, fundamental design question: Should the resolution of conflicting objectives happen at the central server or locally at each client?
@@
Recent work in federated multi-objective supervised learning has explored the server-centric approach to conflict resolution. A straightforward implementation requires each client to transmit gradients for all $M$ objectives to a central server, which then finds a globally consistent update direction (Yang et al., 2023). This design is practically untenable for modern LLMs, as it inflates communication costs by a factor of $M$, resulting in a prohibitive overhead of $\mathcal{O}(CMd)$, where $C$ denotes the number of clients and $d$ the model dimension. Recognizing this bottleneck, other works have proposed using gradient compression before transmission (Askin et al., 2024). However, this remedy introduces significant new overheads: convergence becomes fundamentally limited by the quality of the compression, an extra communication round-trip is required in every step, and the compression itself can be computationally prohibitive for models with billions of parameters. Thus, whether naive or compressed, the server-centric paradigm remains a non-starter for efficient, large-scale LLM alignment.
@@
As our first contribution, we shift the paradigm to a communication-efficient client-centric approach. We introduce Federated In-client Regularized Multi-objective alignment (FIRM), the first framework, to our knowledge, for federated multi-objective LLM alignment. In this framework, each client computes gradients for all $M$ objectives using Proximal Policy Optimization (PPO) (Schulman et al., 2017), and then resolves them locally into a single update direction by solving a Multiple-Gradient Descent Algorithm (MGDA) subproblem (Désidéri, 2012). The central server’s role is simply to aggregate the updated client parameters via FedAvg (McMahan et al., 2017), which reduces communication costs to a practical $\mathcal{O}(Cd)$. However, in deploying this method, we identified a critical instability inherent to naive client-side MGDA resolution, which we term multi-objective disagreement drift. Because the solution to the local MGDA subproblem is highly sensitive to input gradients, minor stochastic variations across clients cause their computed update directions to diverge significantly. When aggregated at the server, these conflicting updates impede learning and prevent effective alignment (see Section 3 and Remark 3 in Section 4.2).
@@ 2 Problem Formulation
Finding Pareto-optimal policies in non-convex settings such as LLM alignment is NP-hard. Thus, we instead target Pareto stationarity, a first-order necessary condition for Pareto optimality (Désidéri, 2012; Zhou et al., 2024).
@@ 3 The FIRM Algorithm for LLM Alignment
In this section, we present our proposed algorithm, Federated In-client Regularized Multi-objective alignment (FIRM). FIRM is designed for practical LLM alignment and proceeds in communication rounds, with each client performing multiple local updates before aggregation. The underlying RL algorithm is a multi-objective adaptation of Proximal Policy Optimization (PPO) (Schulman et al., 2017; Zhou et al., 2024), a standard for LLM fine-tuning. For efficiency, clients fine-tune the LLM using LoRA, meaning only the low-rank adapter matrices are trained and communicated.
@@
We now explain why both terms $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are needed in Eq. (1). The first term, $\mathcal{T}_{1}$, corresponds to the MGDA (Désidéri, 2012). MGDA seeks a convex combination of objective gradients that defines a common descent direction. In a federated setup, however, using only $\mathcal{T}_{1}$ is not sufficient. Without $\mathcal{T}_{2}$, client updates can drift apart. To see this, consider rewriting Eq. (1) as
@@ 4.1 Theoretical Setup and Assumptions
To enable a tractable convergence analysis, we analyze a variant of FIRM, which we refer to as Theoretical-FIRM (TFIRM), where the PPO update is replaced by a foundational actor-critic (AC) framework. This substitution allows us to isolate the core federated and multi-objective challenges. In this AC setting, each client’s local update proceeds in two steps. First, in an inner loop, the critic employs linear function approximation to update its $M$ value function estimates, $\{V_{w}^{j}\}_{j=1}^{M}$, via TD learning on mini-batches of trajectory data. Subsequently, the actor uses these estimates to form the TD-error as an advantage approximation for each objective $j\in[M]$, and then computes the per-objective gradients using the policy gradient theorem (Sutton et al., 1999). Beyond local gradient computation, the federated optimization protocol, including server aggregation and our regularized MGDA solver, remains identical to Algorithm 1. Our theoretical framework is a novel extension of the multi-objective AC algorithm of Zhou et al. (2024) to the federated setting, with the regularized MGDA solver ensuring robust consensus on the descent direction. The complete algorithm is provided in Appendix C (See Algorithm 2), while in the remainder of this section we carry out our analysis under the following standard assumptions.
@@
Assumption 1 imposes standard regularity conditions (Zhou et al., 2024). Condition (a) requires differentiability, which is essential for applying policy gradient methods. Condition (b) ensures ergodicity of the Markov process, yielding a unique stationary distribution under any policy and thereby guaranteeing that long-term objectives are well defined. Finally, condition (c) assumes bounded rewards, a standard requirement to ensure that the value functions remain bounded.
@@
Assumption 2 is standard in the analysis of linear temporal-difference learning (Tsitsiklis and Van Roy, 1999; Xu et al., 2020a, b; Qiu et al., 2021; Zhou et al., 2024). The conditions collectively ensure that the projected Bellman equation is well-posed for any policy $\pi_{\theta}$. Specifically, the negative definiteness of the matrix $\mathbf{A}_{\pi_{\theta}}$, combined with bounded features and rewards, guarantees that the optimal weight vector $\mathbf{w}_{\theta}^{*,j}$ that approximates the value function is unique and uniformly bounded across all policies.
@@
This assumption outlines standard smoothness and boundedness conditions essential for our analysis. The bounded score function (a), a property of common policies like softmax, prevents arbitrarily large gradient updates. The Lipschitz condition on the objective gradients (b) ensures a sufficiently smooth optimization landscape, which is a cornerstone for guaranteeing the convergence of gradient-ascent algorithms to a stationary point (Xu et al., 2020a).
@@ 4.2 Results and convergence analysis
Theorem 1 provides the first convergence guarantee for a federated multi-objective actor-critic algorithm, ensuring that TFIRM converges to a Pareto stationary point. The overall convergence rate, dominated by the optimization error, is $\tilde{\mathcal{O}}(1/T)$. This matches the rate for centralized multi-objective actor-critic methods (Zhou et al., 2024), demonstrating that our federated approach achieves a comparable asymptotic performance. Furthermore, the variance term, $\mathcal{O}(1/(CB))$, exhibits the expected linear speedup with respect to the number of clients $C$ and batch size $B$, consistent with standard analyses in federated learning (Zhang et al., 2024).
@@
The most significant insight from our analysis lies in the characterization of the client drift. The total drift consists of two components. The first, $\mathcal{O}(\alpha^{2}K^{2})$, is the classical drift from local updates, which is well-understood in federated optimization (Zhou et al., 2024; Zhang et al., 2024). The second, and the core of our theoretical contribution, is a novel error term we identify as the multi-objective disagreement drift, given by $\mathcal{O}\!\left(\frac{\sqrt{M^{3}}}{\beta\sqrt{B}}\alpha K\right)$. This term precisely captures the error arising from clients solving the MGDA problem locally with stochastic gradients. Our bound reveals two key insights: (i) the challenge of reaching consensus grows with the number of objectives $M$, and (ii) this drift can be explicitly controlled by the product of the regularization $\beta$ and the batch size $B$.
Comparing to Askin et al. (2024)
While our work is the first in the RL setting, the most relevant methodological comparison is with the federated multi-objective supervised learning framework of Askin et al. (2024). Their server-centric approach tackles client disagreement by having clients communicate compressed gradients to the server, which then solves a single MGDA problem and broadcasts the solution vector $\lambda$ back. This design for supervised learning introduces an error term of $\mathcal{O}(qM)$ into the convergence upper bound, where the factor $q$ depends on the quality of the gradient compression performed at each client. This approach, therefore, introduces significant overhead. A new error source: the convergence is fundamentally limited by the quality of the gradient compression, as represented by the $q$ term. Communication cost: it requires an extra round-trip communication in every training step (gradients up, $\lambda$ down). Computational cost: it relies on a dimensionality reduction step on the gradients at each client, which can be computationally prohibitive for large models. In contrast, our approach removes the need for costly server-side coordination and its associated trade-offs. By introducing a lightweight $\ell_{2}$ regularization to each client’s local MGDA objective, we provably control the multi-objective disagreement drift.
@@
Our upper bound differs from Askin et al. (2024) by an additional $\sqrt{M}$ factor. This may stem from the fact that client drift in RL is more severe than in supervised learning, as local policy updates affect both the local model parameters and the local induced data distribution. Understanding whether this dependence is tight remains an interesting open problem left for future work.
@@
The critic approximation error $\varepsilon_{\text{critic}}$ in our bound can be controlled using standard results from temporal-difference learning. In particular, analyses of Markovian mini-batch TD show that this error can be driven arbitrarily close to zero with sufficient data. From Theorem 1 of Xu et al. (2020a), we obtain $\mathbb{E}[\|w-w^{*}\|^{2}]\leq\epsilon$ with sample complexity $\mathcal{O}(\epsilon^{-1}\log(1/\epsilon))$. Thus, in our convergence analysis, the critic error can be treated as a controllable parameter (see Lemma 3 in Appendix D).
@@ 5 Numerical Experiments
We fine-tune TinyLlama-1.1B-Chat-v1.0  [^1] The model choice reflects realistic edge-device constraints in federated learning, where resources are limited. (Zhang and contributors, 2023) using LoRA adapters, ensuring only adapter weights are trained and communicated. Our alignment task uses $M=2$ objectives, Helpfulness and Harmlessness, with prompts from the Anthropic HH-RLHF (Bai et al., 2022) dataset and rewards evaluated by public reward models [^2] Ray2333/gpt2-large-helpful-reward_model (helpfulness) and Ray2333/gpt2-large-harmless-reward_model (harmlessness) (Yang et al., 2024). Moreover, all reward scores are normalized to the $[0,1]$ range. Each generated response receives a 2-D reward vector $(r_{\mathrm{help}},r_{\mathrm{harm}})$ which drives the PPO updates. Our federated protocol involves $C$ clients over $T$ rounds, where each client performs a fixed number of local PPO-style updates before the server aggregates the LoRA (rank r=16r{=}16 , $\alpha_{\mathrm{lora}}{=}32$, dropout $0.05$) adapters via FedAvg. For the local MGDA subproblem, we solve a regularized Quadratic Program (QP). However, the scale of the gradients, and thus their Gram matrix $G$, can vary dramatically during training. To ensure our regularization $\beta$ has a consistent effect, we first normalize $G$ by its trace, which stabilizes the local optimization (See Appendix A). We set the regularization parameter to $\beta=0.01$. All experiments were conducted on a server equipped with NVIDIA H100 GPUs.
@@
To evaluate FIRM, we compare it against a strong server-centric baseline, which we term FedCMOO-A. This baseline is our adaptation of the federated multi-objective supervised learning algorithm, FedCMOO (Askin et al., 2024), to the LLM alignment setting. In FedCMOO-A, clients send their local multi-objective gradients to the server. The server then solves a single MGDA problem to compute a global trade-off vector $\lambda$ and broadcasts it back to the clients. This architecture avoids the multi-objective disagreement drift by design, as all clients are forced to use the same $\lambda$. To ensure a fair comparison focused purely on the conflict resolution strategy, we do not use gradient compression in FedCMOO (Askin et al., 2024), thereby removing any potential compression error. As shown in Figure 1, FIRM, with 8 clients, achieves comparable or superior performance to FedCMOO-A. One hypothesis for FIRM’s strong performance is the agility of its local updates. FIRM clients can adjust their trade-off vectors $\lambda_{c}$ at every local step, allowing for a more responsive and potentially better balance between objectives. In contrast, FedCMOO-A clients must wait for the server’s global $\lambda$, which can become “stale” or lagged between communication rounds. This lag results in oscillatory weight trajectories, as the server repeatedly overcorrects its global $\lambda$ (See Figure 1(c) and Figure 1(d)).